# references.bib

### v1

@@ Abstract
Since the discovery of ferroelectricity in HfO2 thin films, significant research has focused on Zr-doped HfO2 and solid solution (Hf,Zr)O2 thin films. Functional properties can be further tuned via multilayering, however, this approach has not yet been fully explored in HfO2-ZrO2 films. This work demonstrates ferroelectricity in a 50 nm thick, solution-processed HfO2-ZrO2 multilayer film, marking it as the thickest such film to date exhibiting ferroelectric properties. The multilayer structure was confirmed through transmission electron microscopy (TEM) and energy dispersive x-ray spectroscopy, with high-resolution TEM revealing grain continuity across multiple layers. This finding indicates that a polar phase in the originally paraelectric ZrO2 layer, can be stabilized by the HfO2 layer. The film attains a remanent polarization of 8 µC/cm2 and exhibits accelerated wake-up behavior, attributed to its higher breakdown strength resulting from the incorporation of multiple interfaces. These results offer a faster wake-up mechanism for thick ferroelectric hafnia films.
@@ 1 Introduction
In this work, we investigate the stabilization of ferroelectricity in 50 nm multilayer La::::HfO2-ZrO2 thin films, which represent the thickest multilayer reported to date. We prepared pure ZrO2, La::::HfO2, and multilayer La::::HfO2-ZrO2 films using a CSD process described in our previous publication.[Mandal2024] For clarity in the following discussion, the thin-film samples of La::::HfO2, ZrO2, and La::::HfO2-ZrO2 multilayers are referred to as HO, ZO, and HZO, respectively. It is observed that although the pure ZrO2 film is paraelectric, the combination with the polar HfO2 layer induces ferroelectricity in ZrO2 layer of the HZO film. High-resolution transmission electron microscopy (HR-TEM) confirms that the polar phase extends through multiple layers. The HZO film exhibits low leakage current ferroelectric properties with a remanent polarization (Pr) of 8 µC/cm2, which is comparable to conventional HO films. We further show that the multilayered HZO films can sustain higher electric fields than the HO film alone, resulting in significant reduction of wake-up cycles. The number of cycles required for polarization saturation in HZO decreased tenfold (from 10,000 to 1,000) compared to standard HO films.
@@ 3 Results and discussion
Figure 1(a) presents a structural schematic of the multilayer HZO film, which consists of ten 5 nm-thick layers of alternately deposited La::::HfO2 and ZrO2, starting with La:HfO2 and ending with ZrO2. The bright-field STEM image in Figure 1(b) reveals the expected multilayered stack, indicated by the contrast variation between the La:HfO2 and ZrO2 layers, with two Pt-based electrodes enclosing the stack. EDS spectrum (see Supporting Information) to generate the line profile in Figure 1(c) was extracted from the area outlined by the white box in Figure 1(b). The oscillating behavior of the Hf M and Zr K EDS line profiles in Figure 1(c) further validates the multilayered structure of the film, with the Ti K line indicating where the film terminates. The grazing incident X-ray diffraction (GIXRD) patterns of the HO, ZO, and HZO films are illustrated in Figure 1(d). The patterns show the characteristic features typically observed in HfO2 and ZrO2 thin films, i.e., a pronounced peak at 30.6 ° and a weaker peak at 35.5 °, indicative of the orthorhombic polar planes (111) and (002), respectively. However, note that the structural similarities between the polar orthorhombic and non-polar tetragonal phases lead to significant overlap of their reflections, making it challenging to distinguish them in laboratory XRD experiment. Additionally, a broad shoulder peak between $\sim$31° and 31.5° is also present, which may correspond to a monoclinic phase (P21/c).
@@
The current density-electric field and polarization-electric field hysteresis loops after wake-up cycling are shown in Figure 2. The ZO film does not show any switching current and is therefore paraelectric. The HO and multilayered HZO films exhibit prominent switching current peaks. Both films have a positive remanent polarization (Pr) of 8 µC/cm2, with coercive field (Ec) of 1.2 MV cm-1 and 1.5 MV cm-1 for HZO and HO films, respectively. If the ZrO2 layers were purely dielectric, they would have acted as dead layers, reducing the effective electric field across the film. This, in turn, would result in a decrease in remanent polarization and an increase in the coercive field[Zhang2013], which is not observed in this case. Therefore, it can be concluded that HZO and HO samples have effective ferroelectric layers of identical thicknesses. Interesting to note is that when multilayers are prepared with ZrO2 as an initial layer, ferroelectric switching is hardly present (see Supporting Information). Similar result is reported by van Gent et al.[Johanna2024] in epitaxial rhombohedral Hf1-xZrx-ZrO2 superlattices prepared via pulsed laser deposition, indicating that this effect is intrinsic to HfO2-ZrO2 multilayers, regardless of the fabrication method, type of ferroelectric phase or microstructure.
@@
While the conventional HO and multilayer HZO films show comparable ferroelectric properties, an important difference was observed during the wake-up cycling. Figure 4(a) illustrates the increase of remanent polarization 2Pr with respect to the number of wake-up cycles. Two distinct sets of wake-up cycles were conducted on the HZO and HO films. The first set, referred to as procedure 1, consisted of a 1000-cycle wake-up at an electric field of approximately 3 MV cm-1, exceeding the coercive field of typical HfO2 films. The second set, procedure 2, involved 7000 cycles, performed as seven consecutive 1000-cycle wake-ups, with the electric field gradually increasing from about 0.8 MV cm-1 to around 3 MV cm-1. The HO film was unable to withstand immediate application of the fields above Ec, making wake-up cycling with procedure 1 impossible. HZO film, on the other hand, revealed an escalated wake-up behaviour upon treating with procedure 1. The pristine loops of HZO (from procedure 1) and HO (from procedure 2) films are presented in 4(b), where HZO shows a propeller-shaped hysteresis loop with non-zero remanent polarization, very similar to HfxZr1-xO2 films reported in previous studies.[Lu2016, Mller2012] Experimental observations and theoretical modeling have linked that to an electric field-induced phase transition from the tetragonal (T) phase to the orthorhombic o-III phase.[ReyesLillo2014, starschich_ZrO2_2017] Figure 4(c) presents the hysteresis loops after 1000 cycles, showing that HZO reached a maximum 2Pr of 18 µC/cm2. In comparison, the HO film required 7000 cycles to achieve comparable 2Pr (see Figure 4(d)). When using procedure 2, HZO showed nearly identical wake-up behavior to the HO sample (see Figure 4(a)).
@@
Wake-up field cycling is typically performed in hafnia-based ferroelectrics to improve ferroelectric response through redistribution of defects, typically oxygen vacancies.[Pesic2016, Jiang2020] However, these defects, often located at grain boundaries, interfaces or domain walls, can also form conductive filaments upon exposure to electric field.[Max2018, Xu2024, Lee2023] Performing wake-up cycles directly at or above the coercive field is therefore detrimental to dielectric properties of conventional HO films. We hypothesize that the reason multilayer HZO films can endure direct application of high fields are multiple La::::HfO2-ZrO2 interfaces, which disrupt the formation of conductive filaments along the film thickness. Note that enhanced cycling endurance from 106 to 109 cycles is observed in epitaxial rhombohedral Hf1-xZrx-ZrO2 superlattices, confirming that multilayering is an effective approach to robust ferroelectric hafnia-based films.[Johanna2024] Beyond hafnia, Sun et al.[Sun2016] achieved a breakdown strength of 4.5 MV cm-1 in BZT-BCT multilayer films, compared to 1 MV cm-1 in single-layer BZT and BCT films of the same thickness. They explained the effect by inhibited propagation and growth of ”electric trees” in the films with engineered interfaces.
@@ 4 Conclusion and outlook
In summary, we demonstrated ferroelectricity in solution-processed 50 nm-thick HZO multilayer film. The ferroelectric properties of the multilayer thin film were investigated through the characterization of compositional profiles, crystalline phases, and electrical measurements. Although the pure ZrO2 films processed were paraelectric in nature, the synthesis of a multilayer structure with HfO2 induced a polar phase in ZrO2 through the La-doped HfO2 layer. The HZO film exhibits leakage-free ferroelectric properties, with a remanent polarization of 8 µC cm-2 and a coercive field of 1.2 MV cm-1. Additionally, as prepared HZO films can endure higher electric fields due to the presence of multiple interfaces, which significantly reduces the wake-up cycles - an improvement by an order of magnitude compared to the conventional films. This study highlights a novel pathway to enhance the wake-up properties of ferroelectric HfO2 and ZrO2-based thin films by introducing interfaces, rather than relying solely on solid solutions. This approach can be extended to the growth of thicker multilayer films for piezoelectric applications, as the improved wake-up properties and reduced early breakdown during extended cycling help address the challenges typically associated with thicker films.
@@ Acknowledgments
The raw data generated in this study have been deposited in the Zenodo repository under xxx

### v2

@@ Abstract
Since the discovery of ferroelectricity in HfO2 thin films, significant research has focused on Zr-doped HfO2 and solid solution (Hf,Zr)O2 thin films. Functional properties can be further tuned via multilayering, however, this approach has not yet been fully explored in HfO2-ZrO2 films. This work demonstrates ferroelectricity in a 50 nm thick, solution-processed HfO2-ZrO2 multilayer film, marking it as the thickest multilayer film to date exhibiting ferroelectric properties. The multilayer structure was confirmed through transmission electron microscopy (TEM) and energy dispersive x-ray spectroscopy, with high-resolution TEM revealing grain continuity across multiple layers. This finding indicates that a polar phase in the originally paraelectric ZrO2 layer, can be stabilized by the HfO2 layer. The film attains a remanent polarization of 9 µC/cm2 and exhibits accelerated wake-up behavior, attributed to its higher breakdown strength resulting from the incorporation of multiple interfaces. These results offer a faster wake-up mechanism for thick ferroelectric hafnia films.
@@ 1 Introduction
In this work, we examine the stabilization of ferroelectricity and the wake-up behaviour in 50 nm multilayer La::::HfO2-ZrO2 thin films, representing the thickest multilayer thin films studied to date. While there have been studies on thick Hf0.5Zr0.5O2 solid-solution films,[starschich_ZrO2_2017, Schenk2019] they have not specifically focused on multilayer structures. We prepared pure ZrO2, La::::HfO2, and multilayer La::::HfO2-ZrO2 films using a CSD process described in our previous publication.[Mandal2024] La doping for HfO2 layer was chosen based on density functional theory predictions, which suggested that La’s larger ionic radius and lower electronegativity promote stabilization of the ferroelectric orthorhombic Pca21 phase.[Batra2017, Materlik2018] Experimental findings further validated the predictions, demonstrating a wide doping range and reduced polarization relaxation.[Schroeder2018] For clarity in the following discussion, the thin-film samples of La::::HfO2, ZrO2, and La::::HfO2-ZrO2 multilayers are referred to as HO, ZO, and HZO, respectively. It is observed that although the pure ZrO2 film is paraelectric, the combination with the polar HfO2 layer induces ferroelectricity in ZrO2 layer of the HZO film. High-resolution transmission electron microscopy (HR-TEM) confirms that the polar phase extends through multiple layers. The HZO film exhibits ferroelectric properties with a remanent polarization (Pr) of 9 µC/cm2, which is minor improvement to the conventional HO films. We further show that the multilayered HZO films can sustain higher electric fields than the HO film alone, resulting in significant reduction of wake-up cycles. The number of cycles required for polarization saturation in HZO decreased tenfold (from 10,000 to 1,000) compared to standard HO films.
@@ 3 Results and discussion
Figure 1(a) presents a structural schematic of the multilayer HZO film, which consists of ten 5 nm-thick layers of alternately deposited La::::HfO2 and ZrO2, starting with La:HfO2 and ending with ZrO2. The bright-field STEM image in Figure 1(b) reveals the expected multilayered stack, indicated by the contrast variation between the La:HfO2 and ZrO2 layers, with two Pt-based electrodes enclosing the stack. EDS spectrum (see Supporting Information) to generate the line profile in Figure 1(c) was extracted from the area outlined by the white box in Figure 1(b). The oscillating behavior of the Hf M and Zr K EDS line profiles in Figure 1(c), along with EDS cross-section mapping (see Supporting Information) further validates the multilayered structure of the film, with the Ti K line indicating where the film terminates. The grazing incident X-ray diffraction (GIXRD) patterns of the HO, ZO, and HZO films are illustrated in Figure 1(d). The patterns show the characteristic features typically observed in HfO2 and ZrO2 thin films, i.e., a pronounced peak at 30.6 ° and a weaker peak at 35.5 °, indicative of the orthorhombic polar planes (111) and (002), respectively. However, note that the structural similarities between the polar orthorhombic and non-polar tetragonal phases lead to significant overlap of their reflections, making it challenging to distinguish them in laboratory XRD experiment. Additionally, a broad shoulder peak between $\sim$31° and 31.5° is also present, which may correspond to a monoclinic phase (P21/c).
@@
The current density-electric field and polarization-electric field hysteresis loops after wake-up cycling are shown in Figure 2. The ZO film does not show any switching current and is therefore paraelectric. The HO and multilayered HZO films exhibit prominent switching current peaks. The HO film has a positive remanent polarization (Pr) of 8 µC/cm2, with coercive field (Ec) of 1.5 MV cm-1. HZO has a positive Pr of 9 µC/cm2, with Ec of 1.2 MV cm-1. If the ZrO2 layers were purely dielectric, they would have acted as dead layers, reducing the effective electric field across the film. This, in turn, would result in a decrease in remanent polarization and an increase in the coercive field[Zhang2013], which is not observed in this case. Therefore, it can be concluded that HZO and HO samples have effective ferroelectric layers of almost identical thicknesses. Interesting to note is that when multilayers are prepared with ZrO2 as the initial layer, poor ferroelectric properties are observed in the multilayer (see Supporting Information). This is likely due to the non-polar phase of the initial ZrO2 layer. We can speculate that this could be improved if the polar phase would be stabilized in this first layer by optimizing the processing parameters. Similar result is reported by van Gent et al.[Johanna2024] in epitaxial rhombohedral Hf1-xZrx-ZrO2 superlattices prepared via pulsed laser deposition, indicating that this effect is intrinsic to HfO2-ZrO2 multilayers, regardless of the fabrication method, type of ferroelectric phase or microstructure.
@@
While the conventional HO and multilayer HZO films show comparable ferroelectric properties, an important difference was observed during the wake-up cycling. Figure 4(a) illustrates the increase of remanent polarization 2Pr with respect to the number of wake-up cycles. Two distinct sets of wake-up cycles were conducted on the HZO and HO films. The first set, referred to as procedure 1, consisted of a 1000-cycle wake-up at an electric field of approximately 3 MV cm-1, exceeding the coercive field of typical HfO2 films. The second set, procedure 2, involved 7000 cycles, performed as seven consecutive 1000-cycle wake-ups, with the electric field gradually increasing from about 0.8 MV cm-1 to around 3 MV cm-1. The HO film could not withstand the immediate application of fields exceeding Ec and typically broke down after approximately 10 cycles, making wake-up cycling using procedure 1 impossible. HZO film, on the other hand, revealed an escalated wake-up behaviour upon treating with procedure 1. The pristine loops of HZO (from procedure 1) and HO (from procedure 2) films are presented in 4(b), where HZO shows a propeller-shaped hysteresis loop with non-zero remanent polarization, very similar to HfxZr1-xO2 films reported in previous studies.[Lu2016, Mller2012] Experimental observations and theoretical modeling have linked that to an electric field-induced phase transition from the tetragonal (T) phase to the orthorhombic o-III phase.[ReyesLillo2014, starschich_ZrO2_2017] Furthermore, studies have shown that propeller-shaped hysteresis loops arise due to defect pinning and are also influenced by the evolution of charged domain walls.[Mehmood2020, Zhou2022] Figure 4(c) presents the hysteresis loops after 1000 cycles, showing that HZO reached a maximum 2Pr of 18 µC/cm2. In comparison, the HO film required 7000 cycles to achieve comparable 2Pr (see Figure 4(d)). When using procedure 2, HZO showed nearly identical wake-up behavior to the HO sample (see Figure 4(a)).
@@
Wake-up field cycling is typically performed in hafnia-based ferroelectrics to improve ferroelectric response through redistribution of defects, typically oxygen vacancies.[Pesic2016, Jiang2020] However, these defects, often located at grain boundaries, interfaces or domain walls, can also form conductive filaments upon exposure to electric field.[Max2018, Xu2024, Lee2023] Performing wake-up cycles directly at or above the coercive field is therefore detrimental to dielectric properties of conventional HO films. We hypothesize that the reason multilayer HZO films can endure direct application of high fields are multiple La::::HfO2-ZrO2 interfaces, which disrupt the formation of conductive filaments along the film thickness. Breakdown measurements were performed on both samples (see Supporting Information), and early breakdown was indeed observed in the HO film. Furthermore, endurance improved from 104 cycles in HO films to 105 cycles in HZO films (see Supporting Information). Note that enhanced cycling endurance from 106 to 109 cycles is observed in epitaxial rhombohedral Hf1-xZrx-ZrO2 superlattices, confirming that multilayering is an effective approach to robust ferroelectric hafnia-based films.[Johanna2024] Beyond hafnia, Sun et al.[Sun2016] achieved a breakdown strength of 4.5 MV cm-1 in BZT-BCT multilayer films, compared to 1 MV cm-1 in single-layer BZT and BCT films of the same thickness. They explained the effect by inhibited propagation and growth of ”electric trees” in the films with engineered interfaces. In addition to multilayering, early fatigue in ferroelectric thin films has been reported to be tackled through strategies such as engineering of oxygen vacancies, suppressing defect diffusion, and introducing compositional inhomogeneity. [Bao2023, Kang2024, Yan2024]
@@ 4 Conclusion and outlook
In summary, we demonstrated ferroelectricity in solution-processed 50 nm-thick HZO multilayer film. The ferroelectric properties of the multilayer thin film were investigated through the characterization of compositional profiles, crystalline phases, and electrical measurements. Although the pure ZrO2 films processed were paraelectric in nature, the synthesis of a multilayer structure with HfO2 induced a polar phase in ZrO2 through the La-doped HfO2 layer. The HZO film exhibits ferroelectric properties, with a remanent polarization of 9 µC cm-2 and a coercive field of 1.2 MV cm-1. Additionally, as prepared HZO films can endure higher electric fields due to the presence of multiple interfaces, which significantly reduces the wake-up cycles - an improvement by an order of magnitude compared to the conventional films. This study highlights a novel pathway to enhance the wake-up properties of ferroelectric HfO2 and ZrO2-based thin films by introducing interfaces, rather than relying solely on solid solutions. This approach can be extended to the growth of thicker multilayer films for piezoelectric applications, as the improved wake-up properties and reduced early breakdown during extended cycling help address the challenges typically associated with thicker films.
@@ Acknowledgments
Supporting Information
The Supporting Information is available free of charge on the ACS Publications website and from the author. Polarization hysteresis of ZHO film, EDS and FFT diffractogram of HZO film, leakage and fatigue measurements of HO and HZO films.
@@
The raw data generated in this study have been deposited in the Zenodo repository under https://doi.org/10.5281/zenodo.14893377.